Ramblin Wreck is a firm specializing in engineering components. The firm is publicly traded and is considering the following project:

The project will last 5.00 years with an annual cash flow of $40.00 million. The project will require an initial investment of $140.00 million

The firm must determine the cost of capital to evaluate the project. (The project is within the firm’s normal activities)

Ramblin Wreck, Inc. Financial Data:

STOCK DATA: BOND DATA:
Current Price Per Share $29.00 Current Price Per Bond $926.00
# of Shares 2.00 million # of bonds 20,000.00
Book Value $50 million Annual Coupon Rate 8.00%
Face Value Per Bond $1,000
Maturity 10 years


The risk free rate in the economy is currently 2.00%, while investors have a market risk premium of 8.00%. Ramblin Wreck, Inc. has a beta of 1.50. The tax rate is 36.00%.

What is the NPV of the project? (express in millions, so 1000000 would be 1.00)
Answer #1

Cost of Capital=(weight of equity*cost of equity)+(weight of debt*after tax cost of debt)

Cost of equity=risk free rate+(beta*market risk premium)=2%+(1.5*8%)=14%

Before tax cost of debt needs to be found using RATE function in EXCEL

=RATE(nper,pmt,pv,fv,type)

nper=maturity period=10

pmt=annual coupon=(coupon rate*face value)=(8%*1000)=80

pv=926

fv=1000

=RATE(10,80,-926,1000,0)

RATE=9.16%

Before tax cost of debt=9.16%

After tax cost of debt=Before tax cost of debt*(1-tax rate)=9.16%*(1-36%)=5.86%

Market value of equity=No of shares*share price=2 million*29=58 million

market value of debt=number of bonds*current price of bond=20000*926=18500000=18.52 million

Total value=58+18.52=76.52 million

Weight of equity=Market value of the equity/Total value=58/76.52=75.8%

Weight of debt=Market value of the debt/Total value=18.52/76.52=24.2%

Cost of Capital=(75.8%*14%)+(24.2%*5.86%)=12.03%

Now to find the NPV use NPV function in EXCEL

=NPV(rate, Year1 to year5 cashflows)-Year0 cashflow

=NPV(12.03%,Year1 to year5 cashflows)-140

NPV=$4.08 million

Cost of capital 12.03%
Cashflows
Year0 -140
Year1 40
Year2 40
Year3 40
Year4 40
Year5 40
NPV 4.08